OpenStudy (anonymous):

simplify ^3√27+3^3√16-4^3√54

OpenStudy (anonymous):

\[\sqrt[3]{27} = 3\]\[3*\sqrt[3]{16} = 6\sqrt[3]{2}\]\[4*\sqrt[3]{54} = 12\sqrt[3]{2}\]\[3 + 6\sqrt[3]{2}-12\sqrt[3]{2} = 3-6\sqrt[3]{2}\]

OpenStudy (anonymous):

i didnt know how to do it .. thanx daniel :)

OpenStudy (m):

you're asking similar questions so i think it might be better to slow down and try to figure out the mechanics of it first.